### v1

# Smart Predict–then–Optimize Paradigm for Portfolio Optimization in Real Markets

Abstract
Improvements in return forecast accuracy do not always lead to proportional improvements in portfolio decision quality, especially under realistic trading frictions and constraints. This paper adopts the Smart Predict–then–Optimize (SPO) paradigm for portfolio optimization in real markets, which explicitly aligns the learning objective with downstream portfolio decision quality rather than pointwise prediction accuracy. Within this paradigm, predictive models are trained using an SPO-based surrogate loss that directly reflects the performance of the resulting investment decisions. To preserve interpretability and robustness, we employ linear predictors built on return-based and technical-indicator features and integrate them with portfolio optimization models that incorporate transaction costs, turnover control, and regularization. We evaluate the proposed approach on U.S. ETF data (2015–2025) using a rolling-window backtest with monthly rebalancing. Empirical results show that decision-focused training consistently improves risk-adjusted performance over predict–then–optimize baselines and classical optimization benchmarks, and yields strong robustness during adverse market regimes (e.g., the 2020 COVID-19). These findings highlight the practical value of the Smart Predict–then–Optimize paradigm for portfolio optimization in realistic and non-stationary financial environments.
keywords:
Portfolio optimization , Decision-focused learning , Smart predict–then–optimize , Financial machine learning
journal: Expert Systems with Applications
1 Introduction
Portfolio optimization is a central problem in financial decision-making, where investors seek to allocate capital across multiple assets to balance return and risk. With the increasing availability of financial data, machine learning techniques have been widely adopted to support data-driven portfolio construction and asset allocation. Despite these advances, financial markets remain highly complex systems characterized by nonlinearity, noise, and non-stationarity, making robust and reliable investment decision-making challenging in practice. These characteristics expose a fundamental tension between improving predictive accuracy and achieving high-quality portfolio decisions under realistic market frictions.
A large body of prior work focuses on improving return predictability using advanced time-series models, including recurrent neural networks and Transformer-based architectures. While these models have demonstrated strong performance in various forecasting tasks, their practical benefits in financial applications are often limited by weak periodicity, low signal-to-noise ratios, and frequent regime shifts. Moreover, recent evidence suggests that simple and transparent models can achieve competitive predictive performance, calling into question the necessity of highly complex architectures in noisy financial environments. More importantly, improved forecasting accuracy alone does not guarantee superior portfolio performance.
Most learning-based portfolio methods adopt a predict–then–optimize (PtO) paradigm, in which asset returns or related parameters are first estimated and then treated as fixed inputs to a downstream optimization problem. Although intuitive, this decoupled approach suffers from a fundamental mismatch between the training objective and the ultimate decision goal. In particular, small prediction errors can induce large changes in optimal portfolio weights, especially under realistic constraints and trading frictions, leading to suboptimal investment decisions even when predictive accuracy is high.
To address this limitation, the Smart Predict–then–Optimize (SPO) framework and, more broadly, decision-focused learning (DFL) paradigms have been proposed to explicitly realign the learning objective with downstream decision quality. From a paradigm perspective, SPO departs from the conventional predict–then–optimize workflow by embedding the optimization problem into the training loop, thereby training predictive models based on the quality of the resulting decisions rather than prediction error alone.
In this study, we investigate decision-focused learning for portfolio optimization under a realistic rolling-window backtesting framework. We adopt linear predictors for interpretability and robustness, and integrate them with multiple portfolio optimization formulations within a unified SPO-based architecture. Our contributions can be summarized as follows:
  • 1.
    We systematically evaluate the Smart Predict–then–Optimize paradigm for portfolio optimization under realistic trading constraints and transaction costs, and contrast it with conventional predict–then–optimize approaches.
  • 2.
    We propose robust extensions of the SPO framework that explicitly account for prediction uncertainty and demonstrate their effectiveness during adverse market regimes.
  • 3.
    Through extensive empirical studies on U.S. ETF data across multiple market conditions, including crisis and bull-market periods, we show that decision-focused learning consistently improves portfolio decision quality compared to conventional predict–then–optimize baselines.
2 Literature Review
2.1 Portfolio Optimization Modeling
Modern portfolio optimization originates from the mean–variance framework proposed by Markowitz Markowitz [1952], which formulates asset allocation as a trade-off between expected return and risk. Subsequent studies have extended this framework by introducing alternative risk measures, additional constraints, and techniques to address estimation uncertainty, including the Sharpe ratio Sharpe [1998] and coherent risk measures such as Conditional Value-at-Risk (CVaR) Rockafellar and Uryasev [2000]. Robust portfolio optimization further accounts for parameter uncertainty by optimizing portfolio decisions against worst-case scenarios within prescribed uncertainty sets Goldfarb and Iyengar [2003].
From both theoretical and empirical perspectives, Jagannathan and Ma Jagannathan and Ma [2003] demonstrate that portfolio weight constraints can be interpreted as a form of statistical regularization. Such constraints mitigate estimation error and improve out-of-sample risk performance, thereby providing a methodological foundation for incorporating additional constraints and penalty terms in portfolio optimization models.
In particular, transaction costs and portfolio turnover are commonly modeled through penalty terms in the optimization objective to account for realistic trading frictions Boyd et al. [2017].
2.2 Learning-based Approaches
Learning-based approaches to portfolio optimization can be broadly divided into two paradigms according to how learning and decision-making are integrated: reinforcement learning (RL) methods and predict–then–optimize (PtO) methods.
Reinforcement learning approaches formulate portfolio management as a sequential decision-making problem under a Markov decision process, where an agent learns a policy to maximize cumulative rewards through repeated interaction with a financial environment. A variety of RL Moody and Saffell [2001] and deep reinforcement learning (DRL) Jiang et al. [2017] algorithms have been applied in this context, including DQN Pigorsch and Schäfer [2022], PPO Liu et al. [2024], and DDPG Deng et al. [2017]. While these approaches offer high modeling flexibility, they typically rely on carefully designed reward functions and environment simulators, which may lead to unstable training dynamics and limited interpretability in noisy and non-stationary financial markets Bai et al. [2025].
In contrast, predict–then–optimize (PtO) methods decouple the prediction and decision stages by first learning problem parameters, such as asset returns or costs, and then solving a downstream optimization problem to obtain portfolio decisions. This paradigm preserves the structure of classical portfolio optimization models and enables the explicit incorporation of constraints and domain knowledge.
Within the PtO framework, a wide variety of machine learning and neural network models have been extensively investigated for portfolio optimization. In these approaches, learning models are used to estimate key problem parameters, including expected asset returns, risk measures, or other predictive signals, which are subsequently incorporated into classical portfolio optimization formulations. Representative predictive models include linear regression Deng and Min [2013], support vector machine(SVM) Paiva et al. [2019], tree-based method, and deep neural network (DNN) Freitas et al. [2009], such as recurrent neural network (RNN) and long short-term memory (LSTM) Wang et al. [2020] network. Comprehensive reviews and empirical studies  Krauss et al. [2017], Ma et al. [2021]have examined the effectiveness of these models in portfolio-related applications. By integrating data-driven predictions with optimization-based portfolio construction, PtO methods aim to improve investment performance while maintaining the interpretability and structural advantages of traditional portfolio optimization frameworks.
However, stock market prediction remains a challenging time-series forecasting problem, as financial markets are inherently nonlinear, dynamic, noisy, and chaotic, making direct and accurate prediction difficult Deboeck [1994].
Recent advances in recurrent neural networks such as LSTM Lindemann et al. [2021] and Transformer-based models Wang et al. [2024], Zhang et al. [2023], Nie et al. [2023], Liu et al. [2023] have demonstrated strong effectiveness across a wide range of time-series forecasting tasks. Nevertheless, these architectures often rely on inductive biases such as long-range dependency modeling or periodic pattern extraction, which are particularly suitable for structured or seasonal time series. In contrast, financial time series are characterized by weak periodicity, low signal-to-noise ratios, and frequent regime shifts, limiting the practical gains of highly complex architectures in real-world investment settings.
Moreover, the increased architectural complexity of deep recurrent and Transformer-based models raises concerns regarding training stability and interpretability in high-stakes financial decision-making. For example, recent studies such as DLinear Zeng et al. [2023] question the necessity of complex architectures for time-series forecasting, showing that simple linear models can achieve competitive performance.
Nevertheless, even with simplified and more transparent predictive models, the relationship between prediction accuracy and downstream portfolio decision quality remains indirect. Under these circumstances, conventional predict–then–optimize (PtO) approaches that primarily focus on predictive accuracy may fail to yield optimal investment decisions. This limitation motivates the adoption of Smart Predict–then–Optimize (SPO) and other decision-focused learning frameworks, which explicitly align the learning objective with downstream portfolio optimization performance.
2.3 Smart Predict-then-Optimize (SPO)
The distinction between predict–then–optimize (PtO) and Smart Predict–then–Optimize (SPO) was formally introduced by Elmachtoub and Grigas Elmachtoub and Grigas [2022]. While PtO methods focus on minimizing prediction error independently of the downstream optimization problem, SPO explicitly aligns the learning objective with decision quality. In related literature, this paradigm is also referred to as prediction-focused learning and decision-focused learning Mandi et al. [2024].
In a conventional Predict-then-Optimize (PtO) framework, the predictive model is trained by minimizing a surrogate loss function, such as the mean squared error between predicted and realized parameters. The resulting predictions are then treated as fixed inputs to a downstream optimization problem. While this approach is intuitive and computationally convenient, minimizing prediction error does not necessarily lead to high-quality decisions.
This misalignment arises because small prediction errors can result in large changes in the optimal solution of the optimization problem, especially in constrained or ill-conditioned settings. Consequently, improvements in predictive accuracy do not always translate into better downstream decision quality.
To address this issue, the Smart Predict-then-Optimize (SPO) framework proposes to directly incorporate the downstream optimization problem into the learning objective. Instead of optimizing prediction accuracy alone, SPO minimizes a decision-oriented loss, typically defined in terms of decision regret, thereby explicitly aligning the training objective with the quality of the resulting portfolio decisions.
This perspective is especially relevant in financial applications. As discussed earlier, stock return prediction is inherently difficult due to market noise, non-stationarity, and complex dynamics, making prediction errors inevitable. At the same time, financial decision-making places stringent requirements on robustness, stability, and risk control, where small prediction errors may lead to disproportionately large economic losses. Under such conditions, learning paradigms that explicitly optimize decision quality rather than pointwise prediction accuracy are particularly valuable. Recent work has further emphasized the importance of robustness in decision-focused learning by proposing regret-based loss functions that explicitly hedge against adverse decision outcomes Schutte et al. [2024], motivating the adoption of robust SPO-based approaches for portfolio optimization.
Preliminary ideas related to this study were explored in an earlier conference version Wang and Hasuike [2025]. The present work substantially extends this line of research by providing a systematic evaluation and more comprehensive modeling. Building upon this initial investigation, the present study conducts a more systematic evaluation of SPO-based methods in portfolio optimization.
In addition, this study adopts PyEPO Tang and Khalil [2024] to implement the SPO loss. PyEPO offers a convenient and flexible framework for SPO modeling and has been increasingly used in recent working papers and preprints.
3 Methodology
3.1 Smart Predict then Optimize (SPO) based model
In this study, we adopt a Smart Predict–then–Optimize (SPO) framework for portfolio optimization, where prediction and decision-making are tightly coupled through an optimization layer. Due to the computational and implementation complexity associated with backpropagation through optimization problems, we focus on tractable yet expressive portfolio formulations within a unified decision-focused learning architecture.
To preserve model interpretability and adhere to the principle of Occam’s razor, we consistently employ linear models as predictors across all experimental settings. Linear predictors provide a transparent mapping from input features to asset return estimates, which facilitates interpretability and contributes to improved robustness of the resulting portfolio decisions.
Figure 1: Fig1.
Figure 1 illustrates the overall predictive–optimization pipeline. Given input features $\bm{x}$, a predictor produces estimated returns $\hat{\bm{r}}$, which are then passed to a parameterized portfolio optimization module to obtain the portfolio weights. The resulting decision is evaluated using the SPO+ loss, enabling end-to-end training by propagating gradients through both the prediction model and the optimization layer.
Within this unified structure, different portfolio optimization models are instantiated by specifying different objective functions and optimization parameters, collectively denoted by $\Theta$. These instantiations reflect different investment objectives and practical trading considerations. In the following, we consider three representative formulations: MaxReturn, MaxReturn with transaction costs, and MaxReturn with $\ell_{2}$ weight regularization.
3.1.1 MaxReturnSPO
We first consider the basic MaxReturn formulation, which serves as the simplest instantiation of the SPO framework and forms the foundation for subsequent extensions.
Given an input feature vector $\bm{x}\in\mathbb{R}^{d}$, a trainable predictor $f_{\theta}(\cdot)$ produces an estimate of the asset return vector:
$\hat{\bm{r}}=f_{\theta}(\bm{x}).$ (1)
Based on the predicted returns, the portfolio decision is obtained by solving a return-maximization problem:
$\hat{\bm{w}}=\arg\max_{\bm{w}\in\mathcal{W}}\hat{\bm{r}}^{\top}\bm{w},$ (2)
where $\mathcal{W}$ denotes the feasible set of portfolio weights.
During training, the realized return vector $\bm{r}$ is assumed to be observable, and the corresponding oracle decision is defined as
$\bm{w}^{\star}=\arg\max_{\bm{w}\in\mathcal{W}}\bm{r}^{\top}\bm{w}.$ (3)
To enable end-to-end learning, we adopt the SPO+ surrogate loss, which provides a convex upper bound on the true decision regret:
$\mathcal{L}_{\mathrm{SPO+}}(\hat{\bm{r}},\bm{r})=\max_{\bm{w}\in\mathcal{W}}(2\hat{\bm{r}}-\bm{r})^{\top}\bm{w}-\bm{r}^{\top}\bm{w}^{\star}.$ (4)
3.1.2 MaxReturn with transaction costs
To account for practical trading frictions, we extend the MaxReturnSPO framework by explicitly incorporating transaction costs into the portfolio optimization objective. Let $\bm{w}_{t-1}$ denote the portfolio held in the previous period. The decision variable is obtained by solving
$\hat{\bm{w}}=\arg\max_{\bm{w}\in\mathcal{W}}\left(\hat{\bm{r}}^{\top}\bm{w}-\gamma\lVert\bm{w}-\bm{w}_{t-1}\rVert_{1}\right),$ (5)
where $\gamma>0$ controls the transaction costs intensity.
Notation. Let $\hat{\bm{r}}\in\mathbb{R}^{n}$ denote the predicted return vector, $\bm{r}\in\mathbb{R}^{n}$ the realized return vector, $\bm{w}\in\mathbb{R}^{n}$ the portfolio weight vector, and $\mathcal{W}$ the feasible set of portfolios. Let $\bm{w}_{t-1}$ denote the portfolio held in the previous period and $\gamma>0$ the transaction-fee coefficient. We define the value function
$f(\hat{\bm{r}})=\max_{\bm{w}\in\mathcal{W}}\left[(\hat{\bm{r}}-\bm{r})^{\top}\bm{w}-\gamma\|\bm{w}-\bm{w}_{t-1}\|_{1}\right].$
[…]
3.1.3 MaxReturn with $\ell_{2}$ Weight Regularization
We observe that the previous formulations may yield highly concentrated portfolios, occasionally allocating most of the capital to a single asset. Such extreme allocations are often undesirable in practice due to their instability and sensitivity to prediction errors. To alleviate this issue, we introduce an $\ell_{2}$ regularization term on portfolio weights, which penalizes large positions and encourages more balanced and stable allocations.
Specifically, the portfolio decision is obtained by solving
$\hat{\bm{w}}=\arg\max_{\bm{w}\in\mathcal{W}}\left(\hat{\bm{r}}^{\top}\bm{w}-\gamma\|\bm{w}-\bm{w}_{t-1}\|_{1}-\lambda\|\bm{w}\|_{2}^{2}\right),$ (6)
where $\hat{\bm{r}}\in\mathbb{R}^{n}$ denotes the predicted return vector, $\bm{w}\in\mathbb{R}^{n}$ the portfolio weight vector, $\mathcal{W}$ the feasible set of portfolios, $\bm{w}_{t-1}$ the portfolio held in the previous period, $\gamma>0$ the transaction costs coefficient, and $\lambda>0$ the regularization parameter controlling the strength of the $\ell_{2}$ penalty.
Since the $\ell_{2}$ regularization term does not depend on the predicted returns, the resulting optimization problem remains affine in $\hat{\bm{r}}$, and gradients can still be propagated through the optimization layer using the SPO+ surrogate.
3.1.4 RobustSPO
While the standard SPO framework optimizes portfolio decisions based on point predictions of asset returns, such predictions are inevitably subject to estimation errors. To improve robustness against adverse prediction perturbations, we further extend the MaxReturnSPO formulation by incorporating robustness into the prediction–decision pipeline.
Specifically, we consider multiplicative perturbations applied to the predicted return vector. Let $\bm{\zeta}\in\mathbb{R}^{n}$ denote a perturbation vector and define the perturbed prediction as
$\tilde{\bm{r}}=\hat{\bm{r}}\circ(1+\bm{\zeta}),$ (7)
where $\circ$ denotes the element-wise product. The perturbation is assumed to lie in the uncertainty set
$\mathcal{U}=\left\{\bm{\zeta}:\|\bm{\zeta}\|_{\infty}\leq\rho\right\},$ (8)
where $\rho>0$ controls the maximum relative deviation of each asset return.
Given a perturbed prediction $\tilde{\bm{r}}$, the corresponding portfolio decision is obtained by solving
$\tilde{\bm{w}}=\arg\max_{\bm{w}\in\mathcal{W}}\tilde{\bm{r}}^{\top}\bm{w}.$ (9)
The robust SPO objective is then formulated as a worst-case regret minimization problem:
[…]
where $\bm{w}^{\star}$ denotes the oracle portfolio under the realized return vector $\bm{r}$.
In practice, the inner maximization over the uncertainty set is approximated via Monte Carlo sampling. For each mini-batch, multiple perturbed return predictions are generated, and the maximum SPO+ loss among these scenarios is used for backpropagation. This robust training procedure encourages the predictor to produce portfolio decisions that are less sensitive to unfavorable prediction errors.
3.2 SoftmaxDFL
Motivated by prior work Zhang et al. [2020] that proposes constructing portfolio weights via a softmax layer and directly optimizing financial performance metrics through a loss function, we also consider an alternative learning-based portfolio allocation approach based on a fully differentiable allocation structure. As a differentiable baseline, we consider a neural allocator that directly maps input features to portfolio weights via a softmax layer, denoted as SoftmaxDFL. Unlike SPO-based methods that predict returns and then solve an explicit optimization problem, SoftmaxDFL learns the allocation policy end-to-end without an explicit optimization layer.
Figure 2: Fig2.
Given input features $\bm{x}$, a linear inferencer $f_{\theta}(\cdot)$ first produces return estimates
$\hat{\bm{r}}=f_{\theta}(\bm{x}).$ (11)
A neural allocator $g_{\theta}(\cdot)$ then maps $\hat{\bm{r}}$ directly to portfolio weights via a softmax layer:
[…]
which ensures nonnegative weights that sum to one. Unlike SPO-based methods, SoftmaxDFL does not solve an explicit optimization problem in the forward pass.
The learned allocation is evaluated based on its realized portfolio performance.
MaxReturn optimizes portfolio performance by minimizing the negative realized return:
$\mathcal{L}_{\mathrm{MaxReturn}}=-\,\bm{r}^{\top}\hat{\bm{w}}$. (13)
MaxSharpe considers risk-adjusted performance by minimizing a differentiable surrogate of the Sharpe ratio:
$\mathcal{L}_{\mathrm{MaxSharpe}}=-\frac{\bm{r}^{\top}\hat{\bm{w}}}{\sqrt{\hat{\bm{w}}^{\top}\Sigma\hat{\bm{w}}}}$, (14)
where $\Sigma$ denotes the return covariance matrix estimated from the training window.
3.3 Baselines
We consider two classical baselines for portfolio optimization.
3.3.1 Max Sharpe Ratio
As a traditional optimization-based benchmark, we adopt the classical mean–variance portfolio model. Given the realized return vector $\bm{r}$ and covariance matrix $\Sigma$, the oracle portfolio is obtained by solving
$\bm{w}^{\star}=\arg\max_{\bm{w}\in\mathcal{W}}\frac{\bm{r}^{\top}\bm{w}}{\sqrt{\bm{w}^{\top}\Sigma\bm{w}}}$ (15)
3.3.2 Predict-then-Optimize (PtO)
As a Predict-then-Optimize (PtO) baseline, we combines a linear return predictor with the classical Markowitz model. Specifically, given input features $\bm{x}$, a linear model $f_{\theta}(\cdot)$ is trained to predict asset returns by minimizing the mean squared error:
$\hat{\bm{r}}=f_{\theta}(\bm{x}),\qquad\min_{\theta}\;\|\hat{\bm{r}}-\bm{r}\|_{2}^{2}.$ (16)
After training the predictor, the predicted returns are treated as fixed inputs to the portfolio optimization stage. The portfolio weights are then obtained by solving the Markowitz problem
$\bm{w}^{\mathrm{PtO}}=\arg\max_{\bm{w}\in\mathcal{W}}\hat{\bm{r}}^{\top}\bm{w}.$ (17)
This baseline explicitly decouples return prediction and portfolio optimization, and serves as a standard comparison to decision-focused learning approaches.
4 Experiment Setup
4.1 Data Description
To evaluate the proposed SPO framework for portfolio optimization, we employ daily historical data of U.S. exchange-traded funds (ETFs) as the experimental dataset from 2015/01/01 to 2025/01/01. Each asset in the investment universe corresponds to a tradable ETF, and the dataset spans multiple consecutive years of trading days, enabling a realistic portfolio backtesting environment.
At each trading day, the model takes return-based and technical indicator features derived from historical prices as inputs to predict the next-period asset returns (or equivalently, a cost vector), which are then fed into a portfolio optimization model to generate portfolio allocation weights. Specifically, the input features include:
(1) log returns, capturing short-term price movements;
(2) simple moving averages (SMA) and price bias, reflecting trend-following and mean-reversion characteristics;
(3) relative strength index (RSI) and MACD differences, capturing momentum and trend dynamics;
(4) Bollinger band width, measuring market volatility;
(5) volume-based indicators, reflecting trading activity and market participation.
These features jointly describe the market state of each asset from the perspectives of returns, trends, momentum, volatility, and trading behavior, providing economically meaningful inputs for the prediction models. This data setting naturally aligns with the decision-focused learning paradigm, where prediction quality is evaluated through downstream portfolio decisions and backtesting performance rather than pointwise forecasting accuracy.
4.2 Backtest Strategy
We evaluate all models using a rolling-window backtesting framework with monthly rebalancing, which reflects a realistic investment process. At each rebalancing month $t$, the prediction model is retrained from scratch using only historical information available prior to $t$.
Figure 3: Fig3.
As illustrated in Figure  3, for each month $t$, we construct a fixed-length training window covering the past 12 months [t12,t3][t-12,\,t-3], followed by a time-ordered validation period [t3,t][t-3,\,t]. The window is rolled forward by one month at each rebalancing step.
Within each rolling window, hyperparameters are selected via time-series validation by Optuna, preserving the temporal order of the data. After hyperparameter selection, the model is trained on the corresponding window and used to generate predicted return vectors, which serve as inputs to the portfolio optimization problem. The resulting portfolio is held over the out-of-sample period [t,t+1][t,\,t+1] until the next rebalancing date.
In all backtesting experiments, a proportional transaction costs of 0.005 is applied to account for realistic trading frictions.
4.3 Model Configuration
Table 1: Model Configuration of SPO models
Parameter Values
Proportional transaction costs $\gamma$ 0.005
$\ell_{2}$ weight regularization $\lambda$ 0.42
Robustness radius $\rho$ {0.01, 0.1}
Training epochs 20 to 40
Learning rate $10^{-4}$ to $5\times 10^{-2}$ (log-uniform)
Batch size 63
Optimizer Adam
Table 2: Model Configuration of SoftmaxDFL models
Parameter Values
Hidden Layer 32
Training epochs 20 to 40
Learning rate $10^{-4}$ to $5\times 10^{-2}$ (log-uniform)
Batch size 63
Optimizer Adam
Table 1 and Table 2 summarizes the model configuration used in our experiments. transaction costs and regularization parameters are fixed across all runs, while the robustness radius is varied as $\rho\in\{0.01,0.1\}$. The learning rate and the number of training epochs are selected from predefined ranges using time-series validation, while the batch size and optimizer are kept constant.
5 Result and Discussion
5.1 Overall Backtest Performance
Figure 4: Fig4.
Table 3: Overall backtest performance (2016–2024).Annualized return and volatility are reported on an annual basis.
Strategy Ret. Vol. Sharpe Sortino MaxDD
Softmax-MaxReturn 4.11 13.75 0.362 0.331 -32.20
Softmax-MaxSharpe 1.32 13.02 0.166 0.152 -38.37
RobustSPO ($\rho=0.01$) 9.67 15.09 0.688 0.660 -27.79
RobustSPO ($\rho=0.1$) 8.09 14.96 0.595 0.582 -29.64
PtO Markowitz 9.00 14.76 0.659 0.624 -30.22
SPO+ ($\gamma=0.005$) 10.54 15.75 0.715 0.699 -28.51
SPO+ ($\gamma=0.005,\ \lambda=0.42$) 7.69 15.74 0.550 0.512 -30.18
SPO+ 14.05 19.09 0.785 0.728 -28.71
MaxSharpe 7.13 13.63 0.574 0.514 -26.33
Table 3 reports the overall backtesting performance of all compared strategies in terms of annualized return, volatility, Sharpe ratio, Sortino ratio, and maximum drawdown.
Overall, decision-focused learning approaches consistently outperform predict–then–optimize baselines and traditional portfolio optimization methods on a risk-adjusted basis.
Since the primary objective of this study is to improve decision quality rather than predictive accuracy alone, the predict–then–optimize (PtO) framework serves as the most relevant baseline for comparison.
Among all strategies, SPO+ achieves the highest annualized return (14.05%) and the best risk-adjusted performance, with a Sharpe ratio of 0.785 and a Sortino ratio of 0.728. Although SPO+ exhibits relatively higher volatility, its superior return compensates for the increased risk, resulting in the most favorable overall performance.
The SPO+ with transaction costs variant also performs competitively, delivering strong risk-adjusted returns while explicitly accounting for trading costs. In contrast, introducing the $\ell_{2}$ weight regularization term leads to a more conservative and diversified allocation, reducing both return and Sharpe ratio while maintaining drawdown levels comparable to other baselines.
Softmax-based strategies fail to outperform other approaches in either return or risk-related metrics. Despite their end-to-end differentiable structure, these models do not exhibit clear advantages over predict–then–optimize or decision-focused methods in the present setting. This performance gap may be attributed to the limited interpretability and training instability of deep neural networks when applied to noisy and non-stationary financial data, which can adversely affect the quality of the induced portfolio decisions.
5.2 Performance during the COVID-19 Market Crisis
Figure 5: Cumulative net asset value (NAV) trajectories of selected strategies during the COVID-19 market turmoil in 2020. The inset zooms in on the early crash period from January to April 2020. During this interval, the RobustSPO ($\rho=0.1$) and SPO+ with Fee strategies produce almost overlapping NAV paths, suggesting that these models arrive at highly similar portfolio decisions when facing severe market stress. This behavior indicates that, under extreme market conditions, the SPO+ with Fee solution is already sufficiently conservative due to binding constraints and transaction cost penalties, such that additional robustness does not alter the optimal allocation.
Table 4: Backtest performance during the COVID-19 period (January 2020–December 2020).
Strategy Ret. Vol. Sharpe Sortino MaxDD
Softmax-MaxReturn -6.73 18.21 -0.292 -0.222 -30.37
Softmax-MaxSharpe -4.72 16.51 -0.210 -0.161 -26.65
RobustSPO ($\rho=0.01$) 17.81 14.87 1.182 0.999 -10.71
RobustSPO ($\rho=0.1$) 46.92 18.61 2.170 2.322 -9.58
PtO Markowitz 20.63 28.98 0.796 0.693 -30.22
SPO+ ($\gamma=0.005$) 46.92 18.61 2.170 2.322 -9.58
SPO+ ($\gamma=0.005,\ \lambda=0.42$) -3.51 22.50 -0.046 -0.037 -28.93
SPO+ 35.48 31.49 1.128 0.956 -27.80
MaxSharpe 12.42 10.44 1.179 0.965 -10.00
Table 4 reports the backtesting performance of all strategies during the COVID-19 period from January 2020 to December 2020, which represents an extreme market regime characterized by elevated volatility and abrupt drawdowns. Using the predict–then–optimize (PtO) Markowitz model as the primary baseline, we highlight metrics that outperform PtO to facilitate a clear comparison in terms of decision quality.
Overall, decision-focused learning methods exhibit substantial advantages over the PtO baseline during this turbulent period. Both RobustSPO variants and SPO+ with transaction costs achieve significantly higher risk-adjusted performance, as reflected by markedly improved Sharpe and Sortino ratios, while simultaneously reducing maximum drawdowns relative to PtO. In particular, RobustSPO ($\rho=0.1$) and SPO+ with Fee demonstrate strong downside protection, with maximum drawdowns below 10%, compared to over 30% for the PtO baseline.
From a return perspective, several decision-focused models deliver higher annualized returns than PtO, indicating that explicitly incorporating the downstream optimization problem into the learning objective does not merely reduce risk but can also enhance profitability under stressed market conditions. At the same time, the volatility levels of these models remain comparable to or lower than those of PtO, suggesting a more favorable risk–return trade-off.
In contrast, Softmax-based end-to-end allocation strategies perform poorly during the COVID-19 period, failing to outperform the PtO baseline across all evaluated metrics. This observation suggests that fully differentiable allocation models without an explicit optimization structure may be more susceptible to estimation noise and training instability in highly non-stationary market environments.
Taken together, these results indicate that the benefits of decision-focused learning are particularly pronounced during adverse market regimes. By aligning the learning objective directly with portfolio optimization outcomes, decision-focused models yield more robust and stable investment decisions than predict–then–optimize approaches when market uncertainty is elevated.
5.3 Performance during 2024 Bull Market
Figure 6: Fig6.
Table 5: Backtest performance during the bull market period.
Strategy Ret. Vol. Sharpe Sortino MaxDD
Softmax-MaxReturn 10.66 11.91 0.914 0.845 -7.19
Softmax-MaxSharpe 8.79 12.09 0.760 0.713 -7.02
RobustSPO ($\rho=0.01$) 9.99 9.90 1.015 0.969 -5.94
RobustSPO ($\rho=0.1$) 9.38 10.59 0.903 0.856 -5.99
PtO Markowitz 15.07 8.51 1.699 1.716 -5.47
SPO+ ($\gamma=0.005$) 7.71 10.79 0.745 0.719 -5.94
SPO+ ($\gamma=0.005,\ \lambda=0.42$) 14.89 12.72 1.160 1.092 -8.83
SPO+ 35.96 15.09 2.120 2.167 -6.31
MaxSharpe 22.26 11.64 1.793 1.870 -7.18
During the bull market period, models that explicitly incorporate turnover penalties or robustness constraints exhibit noticeably lower returns than the PtO Markowitz baseline. This outcome is likely attributable to the more conservative investment behavior induced by these modeling choices, which restrains aggressive reallocation and leverage during sustained upward market trends.
This observation suggests that a substantial portion of PtO’s profitability is driven by its strong exposure to bullish market conditions. However, it is noteworthy that the baseline SPO+ model still outperforms the PtO baseline in several key metrics, including annualized return and risk-adjusted performance.
This result indicates that, even without additional robustness or turnover constraints, decision-focused learning can effectively capture upside opportunities while maintaining a favorable risk–return trade-off. In this sense, SPO+ strikes a balance between aggressiveness and stability, benefiting from bullish market dynamics without fully sacrificing robustness to market uncertainty.
6 Conclusion
In this study, we investigate decision-focused learning for portfolio optimization under a realistic rolling-window backtesting framework. The empirical results demonstrate that aligning the learning objective with downstream portfolio decisions leads to consistently improved decision quality across different market regimes, compared with conventional predict–then–optimize approaches based on the Markowitz framework.
Our analysis highlights a fundamental trade-off between aggressiveness and robustness in portfolio construction. While predict–then–optimize methods, including traditional Markowitz-based strategies, tend to benefit more from sustained bullish trends through stronger exposure to return forecasts, decision-focused models place greater emphasis on stability and downside risk control, resulting in superior performance under adverse market conditions. Notably, the baseline SPO+ model achieves a balanced behavior across market regimes, outperforming traditional PtO baselines in several key metrics while maintaining the ability to capture upside opportunities without fully sacrificing robustness to market uncertainty.
These findings underscore the importance of incorporating decision awareness into learning-based portfolio optimization, particularly in environments characterized by noisy and non-stationary financial data, where purely prediction-driven Markowitz-style approaches may be insufficient to ensure robust decision quality.
7 Future Work
Several directions remain for future investigation.
First, while the present study adopts proportional transaction costs for tractability and interpretability, incorporating richer market microstructure effects, such as nonlinear market impact, slippage, and no-trade regions, may further improve realism. Investigating how these frictions interact with decision-focused learning objectives remains an important extension.
In addition, richer risk-aware formulations in the optimization layer, including CVaR, downside-risk measures, and drawdown-aware constraints, offer promising avenues to align decision-focused learning with institutional risk management requirements.
Finally, the robust SPO framework considered in this study may be further generalized by adopting adaptive uncertainty sets or distributionally robust Blanchet et al. [2022] decision-focused objectives, enabling systematic handling of regime shifts and prediction uncertainty. Broader empirical validation across asset universes and market conditions would also help assess the generality of the proposed approach.